# MUSES: The Multi-Sensor Semantic Perception Dataset for Driving under Uncertainty

### v2

@@ Abstract
Achieving level-5 driving automation in autonomous vehicles necessitates a robust semantic visual perception system capable of parsing data from different sensors across diverse conditions. However, existing semantic perception datasets often lack important non-camera modalities typically used in autonomous vehicles, or they do not exploit such modalities to aid and improve semantic annotations in challenging conditions. To address this, we introduce MUSES, the MUlti-SEnsor Semantic perception dataset for driving in adverse conditions under increased uncertainty. MUSES includes synchronized multimodal recordings with 2D panoptic annotations for 2500 images captured under diverse weather and illumination. The dataset integrates a frame camera, a lidar, a radar, an event camera, and an IMU/GNSS sensor. Our new two-stage panoptic annotation protocol captures both class-level and instance-level uncertainty in the ground truth and enables the novel task of uncertainty-aware panoptic segmentation, along with standard semantic and panoptic segmentation. MUSES proves both effective for training and challenging for evaluating models under diverse visual conditions, and it opens new avenues for research in multimodal and uncertainty-aware dense semantic perception. Our dataset and benchmark are publicly available at https://muses.vision.ee.ethz.ch/.**footnotetext: equal contribution
@@ 1 Introduction
Figure 1: Modalities and annotations of the MUSES dataset. First row left to right: frame camera, lidar, radar; second row left to right: event camera, panoptic annotation, and difficulty annotation. Best viewed on a screen.
@@
As uncertainty quantification of the perception module is crucial for certifiable downstream control [6, 52, 31, 4], our specialized annotation protocol accounts for high uncertainty in the semantic content of images in MUSES. In particular, we extend the uncertainty-aware annotation protocol of ACDC [40] to model instance-level uncertainty, besides the class-level uncertainty introduced in [40]. This allows MUSES to support the novel task of uncertainty-aware panoptic segmentation, which constitutes a generalization of both panoptic segmentation [27] and uncertainty-aware semantic segmentation [40]. Ground-truth class-level and instance-level difficulty maps are used in uncertainty-aware panoptic segmentation to evaluate respective class-level and instance-level confidence maps jointly with the standard, hard panoptic predictions via the novel average uncertainty-aware panoptic quality (AUPQ) metric, which rewards labels and confidence values that are consistent with their human counterparts.
@@ 2 Related Work
Two of the earliest and most influential datasets for autonomous driving are KITTI [24], which provides high-resolution stereo images, lidar point clouds, and IMU/GNSS data, and Cityscapes [19], which offers high-quality pixel-level semantic annotations for RGB images. However, neither KITTI nor Cityscapes include any samples for adverse weather or illumination. To address this limitation, several datasets have been proposed that explicitly incorporate visual hazards [56] in their data collection. These hazards can be caused by natural phenomena, such as rain, snow, fog, or night, or by artificial factors, such as lens flare, motion blur, or sensor noise. Some of these datasets are Oxford RobotCar [29], Oxford Radar RobotCar [3], DENSE [5], CADC [36], and Boreas [9]. These datasets provide multi-sensor recordings from different modalities, such as RGB cameras, lidar scanners, radar sensors, or thermal cameras. However, most of them do not provide fine, pixel-level semantic annotations, which are essential for training semantic segmentation models that are crucial for downstream action tasks of embodied agents [59]. Another set of datasets provides dense semantic annotations for RGB images under adverse conditions. These include the small-scale Foggy Driving[38], Foggy Zurich [37], Dark Zurich [39] and Raincouver [47] and the large-scale ACDC [40], Wilddash2 [57] and BDD100K [53]. Although these datasets focus on diverse weather and illumination conditions, they are limited to vision-only data and are not usually sufficient for reliable perception under adverse conditions. Some recent datasets provide multimodal data with dense semantic annotations for adverse lighting conditions, including the panoptic extension [30] of Waymo Open [44], or even adverse weather conditions [2, 22]. They combine RGB frames with lidar point clouds to provide a more comprehensive representation of the scene but do not include further sensors or specialized labeling for adverse conditions. Another relevant dataset is DSEC-Semantic [45, 23], which adds event cameras to the frame-lidar combination, but it only provides semantic pseudo-labels, not high-quality ground truth.
Our dataset, MUSES, aims to fill the gap between the existing datasets for semantic perception under adverse conditions (see Tab. 1). MUSES provides diverse-condition multimodal data from a normal frame camera, a lidar, a radar, and an event camera, and corresponding high-quality 2D pixel-level panoptic annotations. We extend the uncertainty-aware semantic segmentation task introduced in [40] to panoptic segmentation, by accounting for both class- and instance-level uncertainty in predictions and ground truth. The resulting novel task of uncertainty-aware panoptic segmentation is enabled by a specialized annotation protocol, and it is fundamentally different from previous work [43] that proposed a new evaluation metric for standard panoptic segmentation.
@@ 3.1 Sensor Suite
3.2 Annotation
Along with the processed multi-sensor data, MUSES contains high-quality panoptic labels. Figure 2 illustrates that lidar point clouds provide insufficient information for accurate 3D annotation of objects in adverse conditions. We thus label the 2D frame camera images, due to their high resolution and ease of interpretability for annotators. The RGB images are labeled for panoptic segmentation with 19 classes, following the taxonomy of Cityscapes evaluation classes [19]. Even 2D panoptic labeling is highly challenging in adverse conditions because these images contain regions with indiscernible semantic content: the image may be too dark or blurry to recognize an object, or a rain droplet on the lens may partially occlude the field-of-view. To account for such cases, we asked annotators to follow the decision hierarchy shown in Fig. 4 for panoptic labeling. If the object class is indiscernible, the pixel is labeled as unknown_class (and by extension unknown_instance), otherwise it belongs to either a thing class or a stuff class. Pixels outside the 19 defined classes (e.g., a wheelchair) are given the fallback label other_class and treated similarly to stuff. If the pixel belongs to a thing class, the annotator can either assign it to an instance or label it as unknown_instance. This decision hierarchy results in two levels of uncertainty: class uncertainty and instance uncertainty. Instance uncertainty occurs when we know the class of the object but not its instance membership, e.g., for pixels at the blurry boundary between two adjacent cars.
@@ 3.2 Annotation
The two-stage annotation process allows us to assign ternary difficulty levels to our labels (see last column in Fig. 3). Pixels with valid labels that are fully consistent in H1 and H2 are labeled not_difficult. Pixels with a consistent class label in H1 and H2, but an inconsistent instance label or the unknown_instance label, receive a difficult_instance label. All other pixels, i.e., pixels with an inconsistent class label in H1 and H2 or pixels that are unknown_class, receive a difficult_class label. While irrelevant for standard tasks such as semantic segmentation and panoptic segmentation, these difficulty levels enable a new task explained in Sec. 4: uncertainty-aware panoptic segmentation.
@@ 4 Uncertainty-Aware Panoptic Segmentation
The ternary difficulty levels assigned to each pixel during the annotation of MUSES (see Sec. 3.2) enable a novel task: uncertainty-aware panoptic segmentation. In this task, a panoptic segmentation model is compensated for errors in difficult image regions, if it predicts the difficulty level correctly. To incorporate this idea into evaluation, we introduce the uncertainty-aware panoptic quality (UPQ) metric, an extension of panoptic quality [27].
@@ 4.1 Uncertainty-Aware Panoptic Quality
In addition to the panoptic prediction and ground truth, UPQ takes as input a binary class confidence prediction and a binary instance confidence prediction, which are evaluated against the ternary difficulty maps (see Sec. 3.2). This confidence comparison transforms some pixels in the panoptic prediction to either ANY or VOID, as follows: First, the binary class confidence prediction divides all pixels into class-unconfident (CU) and class-confident (CC) sets. CU pixels are compared with the ground truth difficulty map: If they are indeed class_difficult, they are converted to ANY pixels in the panoptic prediction, otherwise, they are converted to VOID. The CC pixels are instead divided into instance-unconfident (IU) and instance-confident (IC) sets according to the binary instance confidence prediction. IU pixels with a correct class prediction are compared with the difficulty map: If they are difficult_instance or difficult_class, they are converted to ANY pixels, otherwise, they are converted to VOID. Analogously to the standard PQ [27], UPQ is composed of two steps: segment matching and UPQ computation given the matched segments. Compared to PQ, UPQ modifies the first step, by incorporating ANY pixels as wild cards and VOID pixels as wrong predictions. Note that, if a model predicts all pixels as confident, no ANY or VOID pixels are created and UPQ reduces to the standard PQ.
@@
    Ignore all ANY pixels in both prediction and ground truth, then match remaining segments with $\text{IoU}>0.5$. After matching, replace ANY pixels within matched segments with the correct labels of their respective matches in the ground truth.
@@
    Match remaining ANY segments with remaining ground-truth segments if $\text{IoU}>0.5$. For matches, replace the ANY label with the ground truth.
Figure 6 conveys the intuition for this process: by ignoring the ANY pixels in step 1, segments surrounded by accurate confidence predictions are more easily matched. Even though the predicted segment has an IoU of less than 0.5 with the ground-truth one, it is still matched because the confidence level of the surrounding pixels is correctly predicted as difficult_class or difficult_instance (which results in ANY pixels). If the confidence level of the entire instance were correctly classified as difficult, it would still be matched in step 2.
@@ 5 Analysis and Experiments
Figure 7: (a) Percentages of pixels labeled in the two annotation stages for different conditions. (b) Distance distribution of instances across different weather conditions. H1: stage 1, H2: stage 2, “n/a”: instances without any lidar returns.
@@ 5.1 Analysis of Annotations
Fig. 8 summarizes the number of annotated pixels per class in MUSES. Overall, 78.5% of all pixels receive a valid label after stage 2 of the annotation (in H2), an increase of 10.3% compared to H1. Notably, H2 contains roughly 1.5× more instances than H1. The increased level of detail and completeness in H2 supports the design choice of complementing the annotation with multi-sensor information. When comparing the annotation density in H2 for images of different visual conditions in Fig. 7(a), we observe that fog and nighttime constitute the most challenging conditions to annotate, showing the greatest increases in annotation density when incorporating auxiliary data in stage 2 of the annotation. To measure the added difficulty of stage 2 annotations, we train Mask2Former [15] with H2 and evaluate it on stage 1 (H1) and stage 2 (H2) labels. We observe a 11.1% drop in PQ, from 58.0% (H1) to 46.9% (H2), which indicates the substantial increase in the difficulty of the labels that are created thanks to our multimodal annotation. This highlights the challenging nature of ground truth in MUSES compared to the ground truth in 2D semantic perception datasets that lack auxiliary data—such as our additional modalities—for adverse-condition annotations. We provide a quantitative verification of this argument via a cross-dataset comparison in Sec. 5.3. As a result of our two-stage annotation protocol, 24.5% of pixels are labeled as difficult_class, and 1.5% of things pixels are labeled as difficult_instance.
We investigate the range distribution of annotated things instances in MUSES by estimating their distances via the motion-compensated and outlier-filtered lidar point cloud. Figure 7(b) shows that annotated instances in adverse conditions are closer on average. However, those results need to be interpreted cautiously, as the lidar can be heavily affected by adverse conditions, e.g., in fog 43.4% of instances have less than 10 lidar returns and 22.3% have none. This underscores the advantage of 2D image labeling over point cloud labeling for adverse scenes. More statistics for MUSES are provided in Appendix 0.H.
@@ 5.2 Benefit of Additional Modalities for Dense Semantic Perception
In Tab. 3, we show the benefit of the additional non-camera modalities which MUSES offers compared to previous dense semantic driving perception benchmarks [19, 53, 40], by comparing multimodal models to a camera-only model. In particular, we train different Mask2Former-based networks [15] on MUSES for panoptic segmentation, each using different modalities as inputs. For this, we construct simple multimodal networks by employing separate Swin-T [28] backbones for each modality and fusing their outputs with a parallel cross-attention block [8]. For more implementation details, please refer to Appendix 0.E.
The key observation on the results of this comparison is that the quadrimodal network outperforms the baseline camera-only network across all visual conditions. This finding proves the value of our novel multimodal dense semantic perception dataset for developing better models for such pixel-level tasks. In addition, we observe that all three examined bimodal networks, which fuse the camera input with one of the additional modalities respectively, also deliver consistent improvements over the camera-only network across conditions and trail the performance of the quadrimodal network.
Each non-camera modality provides valuable complementary information for semantic perception on top of the camera. However, there is significant potential to improve upon our simple fusion method. Future multimodal dense semantic perception works based on MUSES could develop more specialized fusion architectures and enhance model performance across all domains.
@@ 5.4 Uncertainty-Aware Panoptic Segmentation
For the novel uncertainty-aware panoptic segmentation task introduced in Sec. 4, we present in Tab. 6 two mask-classification baselines and an oracle approach both for the unimodal and the multimodal setting. The first baseline simply predicts 100% class and instance confidence for all pixels. In this case, the AUPQ is equivalent to the standard PQ. The second baseline aims to extract class and instance confidence from trained mask-classification models: we obtain the class confidence by marginalizing over probability-mask pairs, whereas the instance confidence is obtained by dividing the mask class score through the class confidence. A more detailed description and motivation for this procedure is given in Sec. 0.E.3. The oracle approach is a confidence oracle: it uses ground-truth class and instance confidence maps for confidence prediction. Its performance sets an upper bound for the examined models. As shown in Tab. 6, the first baseline outperforms the marginalization approach, meaning that the confidence scores of those models are not well calibrated to the difficulty labels. Furthermore, there is a substantial performance gap between the two baselines and the oracle, indicating that future custom-trained models have great room for improvement in this novel task.
@@
Multi- modal Confidence AUSQ $\uparrow$ AURQ $\uparrow$ AUPQ $\uparrow$
@@
$\times$ Constant 100% 78.3 58.0 58.2 31.3 46.9
$\times$ Marginalization 80.2 53.6 54.3 30.7 44.3
$\times$ Oracle 89.7 83.7 78.0 71.2 75.2
Constant 100% 81.1 65.0 63.5 39.9 53.6
Marginalization 82.6 58.0 58.2 36.0 48.8
Oracle 90.7 87.0 81.4 75.7 79.0

### v3

@@ Abstract
Achieving level-5 driving automation in autonomous vehicles necessitates a robust semantic visual perception system capable of parsing data from different sensors across diverse conditions. However, existing semantic perception datasets often lack important non-camera modalities typically used in autonomous vehicles, or they do not exploit such modalities to aid and improve semantic annotations in challenging conditions. To address this, we introduce MUSES, the MUlti-SEnsor Semantic perception dataset for driving in adverse conditions under increased uncertainty. MUSES includes synchronized multimodal recordings with 2D panoptic annotations for 2500 images captured under diverse weather and illumination. The dataset integrates a frame camera, a lidar, a radar, an event camera, and an IMU/GNSS sensor. Our new two-stage panoptic annotation protocol captures both class-level and instance-level uncertainty in the ground truth and enables the novel task of uncertainty-aware panoptic segmentation, along with standard semantic and panoptic segmentation. MUSES proves both effective for training and challenging for evaluating models under diverse visual conditions, and it opens new avenues for research in multimodal and uncertainty-aware dense semantic perception. Our dataset and benchmark are publicly available at https://muses.vision.ee.ethz.ch/.
@@ 1 Introduction
Figure 1: Annotated scene of MUSES. First row left to right: frame camera, lidar, radar; second row left to right: event camera, panoptic annotation, and difficulty annotation.
@@
As uncertainty quantification of the perception module is crucial for certifiable downstream control [6, 55, 32, 4], our specialized annotation protocol accounts for high aleatoric uncertainty in the semantic content of images in MUSES. In particular, we extend the uncertainty-aware annotation protocol of ACDC [42] to model instance-level uncertainty, besides ACDC’s class-level uncertainty. This allows MUSES to support the novel task of uncertainty-aware panoptic segmentation, which constitutes a generalization of both panoptic segmentation [28] and uncertainty-aware semantic segmentation [42]. Ground-truth class-level and instance-level difficulty maps are used in uncertainty-aware panoptic segmentation to evaluate respective class- and instance-level confidence maps jointly with the standard panoptic predictions via the novel average uncertainty-aware panoptic quality (AUPQ) metric, which rewards labels and confidence values that are consistent with their human counterparts.
@@ 2 Related Work
Two of the earliest and most influential datasets for autonomous driving are KITTI [25], which provides high-resolution stereo images, lidar point clouds, and IMU/GNSS data, and Cityscapes [19], which offers high-quality pixel-level semantic annotations for RGB images. However, neither KITTI nor Cityscapes include any samples for adverse weather or illumination. To address this limitation, several datasets have been proposed that explicitly incorporate visual hazards [59] in their data collection. These hazards can be caused by natural phenomena, such as rain, snow, fog, or night, or by artificial factors, such as lens flare, motion blur, or sensor noise. Some of these datasets are Oxford RobotCar [30], Oxford Radar RobotCar [3], DENSE [5], CADC [38], AWARE [37] and Boreas [9]. These datasets provide multi-sensor recordings from different modalities, such as RGB cameras, lidar scanners, radar sensors, or thermal cameras. However, most of them do not provide fine, pixel-level semantic annotations, which are essential for training semantic segmentation models that are crucial for downstream action tasks of embodied agents [62]. Another set of datasets provides dense semantic annotations for RGB images under adverse conditions. These include the small-scale Foggy Driving[40], Foggy Zurich [39], Dark Zurich [41] and Raincouver [49] and the large-scale ACDC [42], Wilddash2 [60] and BDD100K [56]. Although these datasets focus on diverse weather and illumination conditions, they are limited to vision-only data and are not usually sufficient for reliable perception under adverse conditions. Some recent datasets like InfraParis [23] provide multimodal data with dense semantic annotations, some include adverse lighting conditions like [51] and the panoptic extension [31] of Waymo Open [46], or even adverse weather conditions [2, 22]. They combine RGB frames mainly with lidar point clouds to provide a more comprehensive representation of the scene but do not include further sensors or specialized labeling for adverse conditions. Another relevant dataset is DSEC-Semantic [47, 24], which adds event cameras to the frame-lidar combination, but it only provides semantic pseudo-labels.
Our dataset, MUSES, aims to fill the gap between the existing datasets for semantic perception under adverse conditions (see Tab. 1). MUSES provides diverse-condition multimodal data from a normal frame camera, a lidar, a radar, and an event camera, and corresponding high-quality 2D pixel-level panoptic annotations. We extend the uncertainty-aware semantic segmentation task introduced in [42] to panoptic segmentation, by accounting for both class- and instance-level aleatoric uncertainty in predictions and ground truth. The resulting novel task of uncertainty-aware panoptic segmentation is enabled by a specialized annotation protocol, and it is fundamentally different from previous work [45] that investigates epistemic uncertainty for panoptic segmentation.
@@ 3.2 Annotation
Along with the processed multi-sensor data, MUSES contains high-quality panoptic labels. Figure 2 illustrates that lidar point clouds provide insufficient information for accurate 3D annotation of objects in adverse conditions. We thus label the 2D frame camera images, due to their high resolution and ease of interpretability for annotators. The RGB images are labeled for panoptic segmentation with 19 classes, following the taxonomy of Cityscapes evaluation classes [19]. Even 2D panoptic labeling is highly challenging in adverse conditions because these images contain regions with indiscernible semantic content: the image may be too dark or blurry to recognize an object, or a rain droplet on the lens may partially occlude the field-of-view. To account for such cases, we asked annotators to follow the decision hierarchy shown in Fig. 4 for panoptic labeling. If the object class is indiscernible, the pixel is labeled as unknown_class (and by extension unknown_instance), otherwise it belongs to either a thing class or a stuff class. Pixels outside the 19 defined classes (e.g., a wheelchair) are given the fallback label other_class and treated similarly to stuff. If the pixel belongs to a thing class, the annotator can either assign it to an instance or label it as unknown_instance. This decision hierarchy results in two levels of uncertainty: class and instance uncertainty. Instance uncertainty occurs when we know the class but not its instance membership, e.g., for pixels at the blurry boundary between two adjacent cars.
@@
The two-stage annotation process allows us to assign ternary difficulty levels to our labels (see last column in Fig. 3). Pixels with valid labels that are fully consistent in H1 and H2 are labeled not_difficult. Pixels with a consistent class label in H1 and H2, but an inconsistent instance label or the unknown_instance label, receive a difficult_instance label. All other pixels, i.e., pixels with an inconsistent class label in H1 and H2 or pixels that are unknown_class, receive a difficult_class label.
@@ 4 Uncertainty-Aware Panoptic Segmentation
The ternary difficulty levels assigned to each pixel during the annotation of MUSES (see Sec. 3.2) enable a novel task: uncertainty-aware panoptic segmentation. In this task, a panoptic segmentation model is compensated for errors in difficult image regions if it predicts the difficulty level correctly. To incorporate this idea into evaluation, we introduce the uncertainty-aware panoptic quality (UPQ) metric, an extension of panoptic quality [28].
@@ 4.1 Uncertainty-Aware Panoptic Quality
In addition to the panoptic prediction and ground truth, UPQ takes as input a binary class confidence prediction and a binary instance confidence prediction, which are evaluated against the ternary difficulty maps (see Sec. 3.2). This confidence comparison transforms some pixels in the panoptic prediction to either ANY or VOID, as follows: First, the binary class confidence prediction divides all pixels into class-unconfident (CU) and class-confident (CC) sets. CU pixels are compared with the ground truth difficulty map: If they are indeed class_difficult, they are converted to ANY pixels in the panoptic prediction, otherwise, they are converted to VOID. The CC pixels are instead divided into instance-unconfident (IU) and instance-confident (IC) sets according to the binary instance confidence prediction. IU pixels with a correct class prediction are compared with the difficulty map: If they are difficult_instance or difficult_class, they are converted to ANY pixels, otherwise, they are converted to VOID. Thus, ANY masks pixels that are correctly predicted as uncertain (class or instance level), while VOID masks pixels that are erroneously predicted as uncertain. Note that ANY/VOID masks apply to the predictions, not the ground truth.
Analogously to the standard PQ [28], UPQ is composed of two steps: segment matching and UPQ computation given the matched segments. Compared to PQ, UPQ modifies the first step, by incorporating ANY pixels as wild cards and VOID pixels as wrong predictions. Hence, UPQ forgives errors for difficult pixels when predictions are correctly uncertain but penalizes predicting "easy" pixels as uncertain. Note that, if a model predicts all pixels as confident, no ANY or VOID pixels are created and UPQ reduces to the standard PQ.
@@
    Ignore all ANY pixels in both prediction and ground truth, then match remaining segments with $\text{IoU}>0.5$. After matching, copy the ground truth class/instance labels to the ANY pixels within matched segments.
@@
    Check remaining unmatched ground-truth segment for $>50\%$ overlap with ANY pixels. For matches, replace the ANY label with the ground truth.
Figure 6 conveys the intuition for this process: by ignoring the ANY pixels in step 1, segments surrounded by accurate confidence predictions are more easily matched. Even though the predicted segment has an IoU of less than 0.5 with the ground-truth one, it is still matched because the confidence level of the surrounding pixels is correctly predicted as difficult_class or difficult_instance (which results in ANY pixels). Note that ANY pixels never form segments and thus cannot create FP or FN segments. If the confidence level of the entire instance were correctly classified as difficult, it would still be matched in step 2.
@@ 5 Analysis and Experiments
Figure 7: (left) Percentages of pixels labeled in the two annotation stages for different conditions. (right) Distance distribution of instances across different weather conditions. H1: stage 1, H2: stage 2, “n/a”: instances without any lidar returns.
@@ 5.1 Analysis of Annotations
Fig. 8 summarizes the number of annotated pixels per class in MUSES. Overall, 78.5% of all pixels receive a valid label after stage 2 of the annotation (in H2), an increase of 10.3% compared to H1. Notably, H2 contains roughly 1.5× more instances than H1. The increased level of detail and completeness in H2 supports the design choice of complementing the annotation with multi-sensor information. When comparing the annotation density in H2 for images of different visual conditions in Fig. 7 (left), we observe that fog and nighttime constitute the most challenging conditions to annotate, showing the greatest increases in annotation density when incorporating auxiliary data in stage 2 of the annotation. To measure the added difficulty of stage 2 annotations, we train Mask2Former [15] with H2 and evaluate it on stage 1 (H1) and stage 2 (H2) labels. We observe a 11.1% drop in PQ, from 58.0% (H1) to 46.9% (H2), which indicates the substantial increase in the difficulty of the labels that are created thanks to our multimodal annotation. This highlights the challenging nature of ground truth in MUSES compared to the ground truth in 2D semantic perception datasets that lack auxiliary data—such as our additional modalities—for adverse-condition annotations. We provide a quantitative verification of this argument via a cross-dataset comparison in Sec. 5.3. As a result of our two-stage annotation protocol, 24.5% of pixels are labeled as difficult_class, and 1.5% of things pixels are labeled as difficult_instance.
We investigate the range distribution of annotated things instances in MUSES by estimating their distances via the motion-compensated and outlier-filtered lidar point cloud. Figure 7 (right) shows that annotated instances in adverse conditions are closer on average. However, those results need to be interpreted cautiously, as the lidar can be heavily affected by adverse conditions, e.g., in fog 43.4% of instances have less than 10 lidar returns and 22.3% have none. This underscores the advantage of 2D image labeling over point cloud labeling for adverse scenes. More statistics for MUSES are provided in Appendix 0.I.
@@ 5.2 Benefit of Additional Modalities for Dense Semantic Perception
In Tab. 3, we show the benefit of the additional non-camera modalities which MUSES offers compared to previous dense semantic driving perception benchmarks [19, 56, 42], by comparing multimodal models to a camera-only model. In particular, we train different Mask2Former-based networks [15] on MUSES for panoptic segmentation, each using different modalities as inputs. For this, we construct simple multimodal networks by employing separate Swin-T [29] backbones for each modality and fusing their outputs with a parallel cross-attention block [8]. For more implementation details, please refer to Appendix 0.F.
The key observation on the results of this comparison is that the quadrimodal network outperforms the baseline camera-only network across all visual conditions. This finding proves the value of our novel multimodal dense semantic perception dataset for developing better models for such pixel-level tasks. In addition, we observe that all three examined bimodal networks, which fuse the camera input with one of the additional modalities respectively, also deliver consistent improvements over the camera-only network across conditions and trail the performance of the quadrimodal network. Each non-camera modality provides valuable complementary information for semantic perception on top of the camera. However, there is significant potential to improve upon our simple fusion method. Future multimodal dense semantic perception works based on MUSES could develop more specialized fusion architectures and enhance model performance across all domains.
@@ 5.4 Uncertainty-Aware Panoptic Segmentation
Confidence AUSQ $\uparrow$ AURQ $\uparrow$ AUPQ $\uparrow$
@@
Constant 100% 78.3 58.0 58.2 31.3 46.9
Marginalization 80.2 53.6 54.3 30.7 44.3
Oracle 89.7 83.7 78.0 71.2 75.2
For the novel uncertainty-aware panoptic segmentation task introduced in Sec. 4, we present in Tab. 6 two mask-classification baselines and an oracle approach for the unimodal setting. The first baseline simply predicts 100% class and instance confidence for all pixels. In this case, the AUPQ is equivalent to the standard PQ. The second baseline aims to extract class and instance confidence from the trained mask-classification model: we obtain the class confidence by marginalizing over probability-mask pairs, whereas the instance confidence is obtained by dividing the mask class score through the class confidence. A more detailed description and motivation for this procedure is given in Sec. 0.F.3. The oracle approach is a confidence oracle: it uses ground-truth class and instance confidence maps for confidence prediction. Its performance sets an upper bound for the examined models. As shown in Tab. 6, the first baseline outperforms the marginalization approach, meaning that the confidence scores of those models are not well calibrated to the difficulty labels. Furthermore, there is a substantial performance gap between the two baselines and the oracle, indicating that future custom-trained models have great room for improvement in this novel task.